ABC Manufacturing Companys costing system has two direct cost categories: direct materials and direct manufacturing labour.
Answer #1

Solution:

1.)

Direct materials $117,000(7,800 × $15)
Direct labor $585,000(7,800 ×$75)
Variable manufacturing overhead $234,000(7,800 ×$30)
Fixed manufacturing overhead $312,000(7,800 × $40)
Total Standard Manufacturing Costs $1,248,000

2.)

a.)

Direct materials price variance = Purchase quantity × ( actual price - standard price) = 25,000 × ($5.20 - $5)

25,000 ×$0.20 = $5000 U

b.)

Direct materials efficiency variance = standard price × ( actual quantity - standard quantity given actual output)

= $5. × (23100 -(7800 × 3lbs per unit) )

=$5 × (23,100 - 23,400)

=$5 ×300 =$1500 F

c.

Direct labor rate variance = actual hours × ( actual rate - standard rate)

= 40,100 hrs. × ($14.6ph - $15 ph)

=40100 hrs. × 0.40

=$16,040 F

d.)

Direct labor efficiency variance = standard rate × ( actual hours - standard hours given actual output)

=$14.60 ×( (40,100 - (7,800 × 5 hours))

=$14.60 × (40100 - 39000)

=$ 14.60 ×1,100

=$16,060 U

e.)

Total manufacturing overhead spending variance = total actual overhead - total budgeted overhead

= $800,000 - (7800 ×30+7800 ×40)

=$800,000 -( $234,000+$312,000)

=$800,000 - $546,000

=$254,000 U

f.)

Variable overhead efficiency variance = standard variable rate × (actual hours -standard hours)

=$30 × (23,100 - 23,400)

$30 ×300

=$9000 F

g.)

Budgeted production = 40000/5 = 8000

Production volume variance = budgeted overhead rate × (actual units - budgeted units)

=$40 × (7,800 - 8,000)

=$40 × 200

=$8,000 U